OpenStudy (swaqqout_kid1):

The amount of carbon-14 remaining in an organism, A, is equal to the product of the initial amount of carbon-14 in the organism, A0, and one-half to the power of n, where n is the number of half-lives the organism has experienced. If a frog initially contained 4 grams of carbon-14 and the half-life of carbon-14 is 5,730 years, how much carbon-14 remains in the frog after 11,460 years? @sangya21

OpenStudy (swaqqout_kid1):

@sammixboo

OpenStudy (anonymous):

\[A = A_0 * (1/2)^n\] \[A = 4 * (1/2)^2\] \[A_0 = \space intial \space amount\space of \space carbon\] \[n = \space no. \space of\space half-life \space of \space carbon\] n = 11460/5730 = 2 \[A = 4 * (1/2)^2\] \[A = 4 * (1/4)\] \[A = 1 \space grams\]